# Radio observations point to a moderately relativistic outflow in the fast X-ray transient EP241021a

### v1

@@ 6 Discussion
Our observations of EP241021a place it within a growing population of FXRTs that show compelling evidence for relativistic outflows despite the absence of prompt gamma-ray detections (e.g., Cenko et al., 2011; Lipunov et al., 2022; Bright et al., 2024; Perley et al., 2025; Srinivasaragavan et al., 2025; Busmann et al., 2025; Li et al., 2025a). In Fig. 2 we show that most of these gamma-ray dark FXRTs require a soft photon index larger than 1 to explain the lack of gamma-rays, and identify a sub-class of events which implies unusually soft spectra with index $\gtrsim$1.5. EP241021a lies within the bulk of the FXRTs population, in a region that is also occupied by canonical GRBs. Therefore, the lack of gamma-ray signal does not immediately single it out as an outlier of the GRB population. Its moderate redshift ($z=0.75$) indicates that the lack of gamma-rays cannot be attributed to cosmological redshift effects, as might be the case for higher-redshift FXRTs (e.g., EP240315a at $z=4.9$; Liu et al., 2025a; Ricci et al., 2025). Instead, the emission of EP241021a is intrinsically soft, likely reflecting a low spectral peak or a suppressed gamma-ray output. The luminosity of its afterglow signal points to a highly energetic explosion with $E_{\rm K,iso}\gtrsim 10^{53}$ erg (Sect. 5), which requires an unusually low radiative efficiency $\eta_{\gamma}\lesssim 1\%$, at the lower end of the distribution found in typical GRBs (Beniamini & Piran, 2013; Dichiara et al., 2022).
@@
Further support for this interpretation comes from our broadband afterglow modeling. The full radio data set, along with the early optical and X-ray data, can be reproduced by a standard forward shock model describing the interaction of a moderately relativistic jet with a low-density $n\approx 10^{-4}-10^{-1}~{}\mathrm{cm}^{-3}$ uniform medium. Two different scenarios are consistent with our dataset. The former scenario invokes a moderately relativistic ($\Gamma\approx$40) jet seen close to its axis (Top-Hat model in Table 3), in which the gamma-ray emission is likely suppressed by the compactness of the source (Lithwick & Sari, 2001; Matsumoto et al., 2019). In this model, EP241021a differs from classical GRBs and resembles the properties of a dirty fireball (Dermer, 1999), where a baryon-rich outflow fails to produce bright $\gamma$-rays. This could point to a non-standard central engine, such as a magnetar or a relativistic TDE, launching an outflow with atypical properties.
@@
The latter model invokes a highly relativistic structured jet seen off-axis (Gaussian model in Table 3). In this case, our line of sight is still dominated by the emission produced by ejecta expanding with moderate velocities. However, a high Lorentz factor ($\Gamma\gtrsim$130) jet exists and is not detected due to viewing angle effects. In this model, EP241021a could very well be a standard GRB “orphan" of its prompt emission (Rhoads, 1999). Within this off-axis framework, the origin of the initial WXT flare remains unclear. If our line of sight does not intercept the jet energetic core, then we might not be able to detect the traditional prompt emission phase. Thus, the luminous X-ray flare discovered by EP would be produced by the low-$\Gamma$ ejecta. Its onset could mark the break-out of the jet from the progenitor’s stellar envelope (Matzner, 2003; Bromberg et al., 2012), thus favoring the core-collapse of a massive star rather than a TDE. Alternatively, the WXT flare could represent the onset of the afterglow (Kobayashi & Zhang, 2007), produced as the relativistic ejecta begin to decelerate upon interacting with the ambient medium. The early X-ray signal would thus not require a separate emission component, but rather reflects the natural transition from the coasting to the deceleration phase in the jet’s evolution (Gaussian+WXT model, Table 3).
@@
Our findings do not depend sensitively on the origin of the unusual optical rebrightening (Busmann et al., 2025), which substantially diverges from traditional afterglow models. Multiple interpretations were put forward to describe this atypical feature, such as refreshed shock (Busmann et al., 2025), multi-component jets (Gianfagna et al., 2025), or a reactivation of a long-lived central engine (Shu et al., 2025). Our analysis shows that the radio data are mostly consistent with a standard synchrotron radiation from a decelerating, moderately relativistic, wide-angled fireball, and are not substantially affected by any additional emission components probed at higher frequencies.
@@
Figure 8: Radio luminosity lightcurves. EP241021A (red circles) is compared to other FXRTs (e.g., EP240315A, Ricci et al. 2025; EP240414A, Bright et al. 2024) as well as other high-energy transients, including GRBs (circles; Chandra & Frail 2012), relativistic TDEs (squares; Zauderer et al. 2011; Andreoni et al. 2022) and FBOTs (diamonds; Perley et al. 2019; Margutti et al. 2019; Perley et al. 2021).
@@
In Fig. 8 we place the radio counterpart of EP241021a in the broader context of high-energy transients. Its luminosity is substantially higher than FBOTs and low-luminosity GRBs (llGRBs), while the observed timescales differ from the population of merger-driven short GRBs. The radio light curve fits well within the distribution of cosmological GRBs from collapsars (dark grey circles) and relativistic jetted TDEs (squares). Its luminosity and timescale appear indeed similar to AT2022cmc (blue squares; Andreoni et al. 2022; Pasham et al. 2023).
@@
Additional observations of gamma-ray dark FXRTs are essential to distinguish between different models and to understand the physical origin of this new population of high-energy transients. For instance, the total energetics, temporal profile and variability of the early X-ray emission detected by WXT would help clarify whether it is similar to erratic and highly variable prompt emission flares from an on-axis jet or consistent with other emission mechanisms visible from off-axis observers. Important clues might also come from the study of the environment, which would help differentiate between TDEs, core-collapse and other types of stellar explosions (e.g., Fruchter et al., 2006; van Velzen et al., 2021; O’Connor et al., 2025).
@@ Acknowledgements
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author.
@@ Appendix A Gamma-ray and X-ray flux values for Fast X-ray Transients
1 EP240305a $<5.8$ This work $83_{-14}^{+20}$ Liu et al. 2024a
2 EP240315a $118_{-36.8}^{+47.3}$ Svinkin et al. 2024a $53_{-7}^{+10}$ Zhang et al. 2024
3 EP240413a $<4.1$ This work $110\pm 10$ Lian et al. 2024a
4 EP240417a $<5.3$ This work $9.8_{-8}^{+0.8}$ Hu et al. 2024a
5 EP240426b $<4.4$ This work $36\pm 6$ Pan et al. 2024a
6 EP240506a $<6.5$ This work $110_{-20}^{+30}$ Li et al. 2024a
7 EP240617a $<5.2$ This work $350\pm 30$ Zhou et al. 2024a
8 EP240618a $<0.85$ Ravasio & Jonker 2024 $290_{-60}^{+70}$ Sun et al. 2024b
9 EP240625a $<7.4$ This work $29_{-5}^{+7}$ Pan et al. 2024b
10 EP240626a $<7.3$ This work $170_{-50}^{+73}$ Wu et al. 2024a
11 EP240702a $<5.3$ This work $540_{-120}^{+150}$ Chen et al. 2024
12 EP240703a $39.4\pm 5.63$ Frederiks et al. 2024a $250_{-30}^{+40}$ Wang et al. 2024a
13 EP240801a $<4.1$ This work $48\pm 31$ Zhou et al. 2024b
14 EP240802a $124.9_{-40.0}^{+47.0}$ Frederiks et al. 2024c $170_{-30}^{+40}$ Wang et al. 2024b
15 EP240804a $22\pm 9.45$ Frederiks et al. 2024b $61_{-16}^{+26}$ Wang et al. 2024c
16 EP240807a $18.45_{-3.48}^{+2.32}$ Ridnaia et al. 2024 $17_{-6}^{+7}$ Fu et al. 2024
17 EP240816a $<7.8$ This work $30\pm 5$ Yang et al. 2024
18 EP240820a $<4.8$ This work $12_{-5}^{+7}$ Liang et al. 2024a
19 EP240913a $37\pm 0.2$ Fermi GBM Team 2024 $11_{-4}^{+5}$ Li et al. 2024b
20 EP240918b $<6.7$ This work $26_{-12}^{+41}$ Liang et al. 2024b
21 EP240918c $<6.7$ This work $150_{-20}^{+60}$ Liang et al. 2024b
22 EP240919a $4.5\pm 0.5$ Roberts et al. 2024 $70_{-15}^{+18}$ Liang et al. 2024c
23 EP240930a $26\pm 1$ Smith et al. 2024 $380\pm 50$ Tian et al. 2024
24 EP241021a $<3.7$ This work $33_{-16}^{+48}$ Hu et al. 2024c
25 EP241025a $11.8\pm 0.5$ Godwin & Fermi GBM Team 2024 $16\pm 7$ Li et al. 2024c
26 EP241026a $41\pm 2$ Trigg & Fermi GBM Team 2024 $160\pm 30$ Li et al. 2024d
27 EP241026b $<5.2$ This work $12_{-3}^{+4}$ Lian et al. 2024b
28 EP241030a $37.6\pm 0.57$ de Barra et al. 2024 $7.5_{-2.4}^{+3}$ Wu et al. 2024b
29 EP241101a $<3.6$ This work $120_{-40}^{+50}$ Liang et al. 2024d
30 EP241103a $<4.9$ This work $380_{-90}^{+110}$ Zhao et al. 2024
31 EP241104a $32\pm 7.6$ Sonawane et al. 2024 $20_{-11}^{+10}$ Zhou et al. 2024c
32 EP241107a $<4.2$ This work $10\pm 0.5$ Li et al. 2024e
33 EP241113a $<5.2$ This work $55_{-7.6}^{+12.6}$ Liu et al. 2024b
34 EP241115a $15\pm 2$ Svinkin et al. (2024c) $14.2_{-4}^{+6}$ Hu et al. 2024d
35 EP241202b $<4.5$ This work $54_{-16}^{+20}$ Zhou et al. 2024d
36 EP241206a $<6.2$ This work $49.2_{-12.4}^{+11.9}$ Yin et al. 2024b
37 EP241217a $<4.5$ This work $73\pm 27$ Zhou et al. 2024f
38 EP241217b $5.3\pm 0.2$ Ravasio et al. 2024 $119\pm 10$ Zhou et al. 2024e
39 EP250101a $<4.5$ This work $4.4_{-1.7}^{+2.3}$ Liang et al. 2025
40 EP250108a $<2.6$ Ravasio et al. 2025 $4.2_{-0.9}^{+1.2}$ Li et al. 2025b
41 EP250109a $13\pm 1$ Pathak et al. 2025a $250_{-120}^{+460}$ Li et al. 2025c
42 EP250111a $<4.5$ This work $139_{-40}^{+53}$ Zhao et al. 2025
43 EP250125a $<7.5$ This work $180_{-50}^{+70}$ Wu et al. 2025
44 EP250205a $6.2\pm 0.2$ Mukherjee et al. 2025 $42\pm 11$ Liu et al. 2025b
45 EP250207a $<3.5$ This work $29_{-7}^{+9}$ Li et al. 2025d
46 EP250207b $<6$ This work $61_{-25}^{+42}$ Zhou et al. 2025
47 EP250212a $<4.2$ This work $90_{-30}^{+100}$ Li et al. 2025e
48 EP250223a $<4.6$ This work $44_{-11}^{+14}$ Wang et al. 2025
49 EP250225a $<6.2$ This work $4.3_{-1.3}^{+2.1}$ Jiang et al. 2025b
50 EP250226a $32\pm 1$ Pathak et al. 2025b $2.89\pm 0.08$ Jiang et al. 2025c
51 EP250227a $<3.9$ This work $210_{-120}^{+260}$ Wen et al. 2025
52 EP250302a $<4.6$ This work $700_{-160}^{+200}$ Dai et al. 2025
53 EP250304a $<4.3$ This work $53\pm 4$ Zhang et al. 2025
54 EP250321a $<5.3$ This work $173_{-19}^{+21}$ Hu et al. 2025
55 EP250321b $<7.2$ This work $7.6_{-3}^{+5.3}$ Cheng et al. 2025

### v2

@@ 6 Discussion
Our observations of EP241021a place it within a growing population of FXRTs that show compelling evidence for relativistic outflows despite the absence of prompt gamma-ray detections (e.g., S. B. Cenko et al., 2011; V. M. Lipunov et al., 2022; J. S. Bright et al., 2024; D. A. Perley et al., 2025; G. P. Srinivasaragavan et al., 2025; M. Busmann et al., 2025; W. X. Li et al., 2025). In Fig. 2 we show that most of these gamma-ray dark FXRTs require a soft photon index larger than 1 to explain the lack of gamma-rays, and identify a sub-class of events which implies unusually soft spectra with index $\gtrsim$1.5. EP241021a lies within the bulk of the FXRTs population, in a region that is also occupied by canonical GRBs. Therefore, the lack of gamma-ray signal does not immediately single it out as an outlier of the GRB population. Its moderate redshift ($z=0.75$) indicates that the lack of gamma-rays cannot be attributed to cosmological redshift effects, as might be the case for higher-redshift FXRTs (e.g., EP240315a at $z=4.9$; Y. Liu et al., 2025; R. Ricci et al., 2025). Instead, the emission of EP241021a is intrinsically soft, likely reflecting a low spectral peak or a suppressed gamma-ray output. The luminosity of its afterglow signal points to a highly energetic explosion with $E_{\rm K,iso}\gtrsim 10^{53}$ erg (Sect. 5), which requires an unusually low radiative efficiency $\eta_{\gamma}\lesssim 1\%$, at the lower end of the distribution found in typical GRBs (P. Beniamini & T. Piran, 2013; S. Dichiara et al., 2022).
@@
Further support for this interpretation comes from our broadband afterglow modeling. The full radio data set, along with the early optical and X-ray data, can be reproduced by a standard forward shock model describing the interaction of a moderately relativistic jet with a low-density $n\approx 10^{-4}-10^{-1}~\mathrm{cm}^{-3}$ uniform medium. Two different scenarios are consistent with our dataset. The former scenario invokes a moderately relativistic ($\Gamma\approx$40) jet seen close to its axis (Top-Hat model in Table 3), in which the gamma-ray emission is likely suppressed by the compactness of the source (Y. Lithwick & R. Sari, 2001; T. Matsumoto et al., 2019). In this model, EP241021a differs from classical GRBs and resembles the properties of a dirty fireball (C. D. Dermer, 1999), where a baryon-rich outflow fails to produce bright $\gamma$-rays. This could point to a non-standard central engine, such as a magnetar or a relativistic TDE, launching an outflow with atypical properties.
@@
The latter model invokes a highly relativistic structured jet seen off-axis (Gaussian model in Table 3). In this case, our line of sight is still dominated by the emission produced by ejecta expanding with moderate velocities. However, a high Lorentz factor ($\Gamma\gtrsim$130) jet exists and is not detected due to viewing angle effects. In this model, EP241021a could very well be a standard GRB “orphan” of its prompt emission (J. E. Rhoads, 1999). Within this off-axis framework, the origin of the initial WXT flare remains unclear. If our line of sight does not intercept the jet energetic core, then we might not be able to detect the traditional prompt emission phase. Thus, the luminous X-ray flare discovered by EP would be produced by the low-$\Gamma$ ejecta. Its onset could mark the break-out of the jet from the progenitor’s stellar envelope (C. D. Matzner, 2003; O. Bromberg et al., 2012), thus favoring the core-collapse of a massive star rather than a TDE. Alternatively, the WXT flare could represent the onset of the afterglow (S. Kobayashi & B. Zhang, 2007), produced as the relativistic ejecta begin to decelerate upon interacting with the ambient medium. The early X-ray signal would thus not require a separate emission component, but rather reflects the natural transition from the coasting to the deceleration phase in the jet’s evolution (Gaussian+WXT model, Table 3).
@@
Our findings do not depend sensitively on the origin of the unusual optical rebrightening (M. Busmann et al., 2025), which substantially diverges from traditional afterglow models. Multiple interpretations were put forward to describe this atypical feature, such as refreshed shock (M. Busmann et al., 2025), multi-component jets (G. Gianfagna et al., 2025), or a reactivation of a long-lived central engine (X. Shu et al., 2025). Our analysis shows that the radio data are mostly consistent with a standard synchrotron radiation from a decelerating, moderately relativistic, wide-angled fireball, and are not substantially affected by any additional emission components probed at higher frequencies. To test the robustness of our results, we included the optical re-brightening and modeled the afterglow with a two-component Top-Hat jet, where both jets share the same $n$, $\theta_{c}$ and $\theta_{\mathrm{obs}}$ and have independent $E_{K,\mathrm{iso}}$, microphysical parameters and $\Gamma_{0}$. This shifts some of the inferred physical parameters (e.g., isotropic energy, circumburst density and viewing angle), but not the main inference of a wide-angled relativistic fireball (see Appendix).
@@
Figure 8: Radio luminosity lightcurves. EP241021A (red circles) is compared to other FXRTs (e.g., EP240315A, R. Ricci et al. 2025; EP240414A, J. S. Bright et al. 2024) as well as other high-energy transients, including GRBs (circles; P. Chandra & D. A. Frail 2012), relativistic TDEs (squares; B. A. Zauderer et al. 2011; I. Andreoni et al. 2022) and FBOTs (diamonds; D. A. Perley et al. 2019; R. Margutti et al. 2019; D. A. Perley et al. 2021).
@@
In Fig. 8 we place the radio counterpart of EP241021a in the broader context of high-energy transients. Its luminosity is substantially higher than FBOTs and low-luminosity GRBs (llGRBs), while the observed timescales differ from the population of merger-driven short GRBs. The radio light curve fits well within the distribution of cosmological GRBs from collapsars (dark grey circles) and relativistic jetted TDEs (squares). Its luminosity and timescale appear indeed similar to AT2022cmc (blue squares; I. Andreoni et al. 2022; D. R. Pasham et al. 2023).
@@
Additional observations of gamma-ray dark FXRTs are essential to distinguish between different models and to understand the physical origin of this new population of high-energy transients. For instance, the total energetics, temporal profile and variability of the early X-ray emission detected by WXT would help clarify whether it is similar to erratic and highly variable prompt emission flares from an on-axis jet or consistent with other emission mechanisms visible from off-axis observers. Important clues might also come from the study of the environment, which would help differentiate between TDEs, core-collapse and other types of stellar explosions (e.g., A. S. Fruchter et al., 2006; S. van Velzen et al., 2021; B. O’Connor et al., 2025).
@@ Appendix A Gamma-ray and X-ray flux values for Fast X-ray Transients
1 EP240305a <3.8<3.8 This work $83_{-14}^{+20}$ M. J. Liu et al. 2024
2 EP240315a $118_{-36.8}^{+47.3}$ D. Svinkin et al. 2024b $53_{-7}^{+10}$ W. J. Zhang et al. 2024
3 EP240413a <2<2 This work $110\pm 10$ T. Y. Lian et al. 2024b
4 EP240417a <3.5<3.5 This work $9.8_{-8}^{+0.8}$ J. W. Hu et al. 2024d
5 EP240426b <4.0<4.0 This work $36\pm 6$ X. Pan et al. 2024a
6 EP240506a <6.2<6.2 This work $110_{-20}^{+30}$ D. Y. Li et al. 2024c
7 EP240617a <4.1<4.1 This work $350\pm 30$ H. Zhou et al. 2024d
8 EP240618a <0.85<0.85 M. E. Ravasio & P. G. Jonker 2024 $290_{-60}^{+70}$ H. Sun et al. 2024b
9 EP240625a <5.7<5.7 This work $29_{-5}^{+7}$ X. Pan et al. 2024b
10 EP240626a <4.8<4.8 This work $170_{-50}^{+73}$ Q. Y. Wu et al. 2024
11 EP240702a <4.3<4.3 This work $540_{-120}^{+150}$ W. Chen et al. 2024
12 EP240703a $39.4\pm 5.63$ D. Frederiks et al. 2024a $250_{-30}^{+40}$ Y. L. Wang et al. 2024a
13 EP240801a <3.6<3.6 This work $48\pm 31$ H. Zhou et al. 2024e
14 EP240802a $1249_{-40.0}^{+47.0}$ D. Frederiks et al. 2024b $170_{-30}^{+40}$ Y. L. Wang et al. 2024b
15 EP240804a $22\pm 9.45$ D. Frederiks et al. 2024c $61_{-16}^{+26}$ Y. Wang et al. 2024b
16 EP240807a $18.45_{-3.48}^{+2.32}$ A. Ridnaia et al. 2024 $17_{-6}^{+7}$ Y. C. Fu et al. 2024
17 EP240816a <9.3<9.3 This work $30\pm 5$ H. N. Yang et al. 2024
18 EP240820a <3.1<3.1 This work $12_{-5}^{+7}$ Y. F. Liang et al. 2024b
19 EP240913a $37\pm 0.2$ Fermi GBM Team 2024 $11_{-4}^{+5}$ D. Y. Li et al. 2024d
20 EP240918b <3.7<3.7 This work $26_{-12}^{+41}$ Y. F. Liang et al. 2024c
21 EP240918c <3.8<3.8 This work $150_{-20}^{+60}$ Y. F. Liang et al. 2024c
22 EP240919a $4.5\pm 0.5$ O. J. Roberts et al. 2024 $70_{-15}^{+18}$ Y. F. Liang et al. 2024d
23 EP240930a $26\pm 1$ J. Smith et al. 2024 $380\pm 50$ X. Tian et al. 2024
24 EP241021a <3.2<3.2 This work $33_{-16}^{+48}$ J. W. Hu et al. 2024c
25 EP241025a $11.8\pm 0.5$ M. Godwin & Fermi GBM Team 2024 $16\pm 7$ D. Y. Li et al. 2024b
26 EP241026a $41\pm 2$ A. Trigg & Fermi GBM Team 2024 $160\pm 30$ D. Y. Li et al. 2024a
27 EP241026b <4.8<4.8 This work $12_{-3}^{+4}$ T. Y. Lian et al. 2024a
28 EP241030a $37.6\pm 0.57$ C. de Barra et al. 2024 $7.5_{-2.4}^{+3}$ H. Z. Wu et al. 2024
29 EP241101a <3.2<3.2 This work $120_{-40}^{+50}$ Y. F. Liang et al. 2024a
30 EP241103a <5.5<5.5 This work $380_{-90}^{+110}$ T. Zhao et al. 2024
31 EP241104a $32\pm 7.6$ R. Sonawane et al. 2024 $20_{-11}^{+10}$ H. Zhou et al. 2024a
32 EP241107a <3.8<3.8 This work $10\pm 0.5$ R. Z. Li et al. 2024
33 EP241113a <4.0<4.0 This work $55_{-7.6}^{+12.6}$ Z. Y. Liu et al. 2024
34 EP241115a $15\pm 2$ D. Svinkin et al. (2024c) $14.2_{-4}^{+6}$ J. W. Hu et al. 2024a
35 EP241202b <4.5<4.5 This work $54_{-16}^{+20}$ C. Zhou et al. 2024
36 EP241206a <4.8<4.8 This work $49.2_{-12.4}^{+11.9}$ Y. H. I. Yin et al. 2024
37 EP241217a <4.1<4.1 This work $73\pm 27$ H. Zhou et al. 2024c
38 EP241217b $5.3\pm 0.2$ M. E. Ravasio et al. 2024 $119\pm 10$ H. Zhou et al. 2024b
39 EP250101a <8.0<8.0 This work $4.4_{-1.7}^{+2.3}$ Y. F. Liang et al. 2025
40 EP250108a <2.6<2.6 M. E. Ravasio et al. 2025 $4.2_{-0.9}^{+1.2}$ R. Z. Li et al. 2025a
41 EP250109a $13\pm 1$ U. Pathak et al. 2025a $250_{-120}^{+460}$ R. Z. Li et al. 2025b
42 EP250111a <8.0<8.0 This work $139_{-40}^{+53}$ T. Zhao et al. 2025
43 EP250125a <5.1<5.1 This work $180_{-50}^{+70}$ Q. Y. Wu et al. 2025
44 EP250205a $6.2\pm 0.2$ O. Mukherjee et al. 2025 $42\pm 11$ Z. Y. Liu et al. 2025
45 EP250207a <2.2<2.2 This work $29_{-7}^{+9}$ D. Y. Li et al. 2025b
46 EP250207b <5.4<5.4 This work $61_{-25}^{+42}$ X. Y. Zhou et al. 2025
47 EP250212a <3.8<3.8 This work $90_{-30}^{+100}$ D. Y. Li et al. 2025a
48 EP250223a <4.7<4.7 This work $44_{-11}^{+14}$ Y. Wang et al. 2025
49 EP250225a <3.7<3.7 This work $4.3_{-1.3}^{+2.1}$ S. Q. Jiang et al. 2025a
50 EP250226a $32\pm 1$ U. Pathak et al. 2025b $980$ S. Q. Jiang et al. 2025b
51 EP250227a <2.9<2.9 This work $210_{-120}^{+260}$ W. F. Wen et al. 2025
52 EP250302a <2.6<2.6 This work $700_{-160}^{+200}$ C. Y. Dai et al. 2025
53 EP250304a <3.9<3.9 This work $53\pm 4$ Y. J. Zhang et al. 2025
54 EP250321a <3.7<3.7 This work $173_{-19}^{+21}$ D. F. Hu et al. 2025
55 EP250321b <5.7<5.7 This work $7.6_{-3}^{+5.3}$ H. Q. Cheng et al. 2025
Appendix B Two-component TopHat Jet Fit (Including Optical Bump)
Figure 9: Broadband modeling including optical re-brightening results. Multi-band light curves fitted with a two-component uniform (Top-Hat) jet model. X-ray (Swift/XRT; 1 keV), optical/NIR ($r$, $z$, $J$; M. Busmann et al. 2025), and 15.5 GHz radio (AMI-LA; F. Carotenuto et al. 2024) data are shown. Downward arrows mark $3\sigma$ upper limits; open symbols indicate points excluded in the fit. Solid lines show the best-fitting two-component jet model, with contributions from component 1(dashed) and second component(dotted) indicated separately. Band-dependent scaling factors are applied for clarity.
Figure 10: Broadband modeling including optical re-brightening results. Posterior distributions of the fitted parameters for the uniform (Top-Hat) jet model.